OpenStudy (anonymous):

When computing the nth derivative of atan(x^3/6), I've tried approaching the problem so that if atan(x) = 1/(1+x^2), then atan(x^3/6) = 1/(1+(x^3/6)^2). That's ok for the first few, but by the time you get to say, the ninth or tenth derivative, things get hairy in a hurry. Should I just be looking for the coefficient of the 9th term in the MacLaurin expansion of 1/(1+x^2)?

OpenStudy (anonymous):

is this arctangent?

OpenStudy (anonymous):

this statement \[\arctan(x)=\frac{1}{1+x^2}\] cannot be right

OpenStudy (anonymous):

that is the derivative

OpenStudy (anonymous):

Ah yes, sorry that's what I meant to say. That I used that as the first derivative, so that \[\frac{ d }{ dx }\tan^{-1}(\frac{x^{6}}{3}) = (1+x^{2})^{-1}\]

OpenStudy (anonymous):

you need the chain rule for this derivative

OpenStudy (anonymous):

alternatively if what you are looking for is the mclaurin series write the series for \(\arctan(x)\) and replace \(x\) by \(\frac{x^6}{3}\)

OpenStudy (anonymous):

So, I have \[\frac{d}{dx}\tan^{-1}(\frac{x^{6}}{3}) = (1+(\frac{x^{6}}{3})^{2})^{-1}\] Where I think I'm freaking out is in taking that out to the 9th derivative to find \[f^{9}(0) \]

OpenStudy (anonymous):

you are doing too much work and in any case that is not right

OpenStudy (anonymous):

the derivative requires the chain rule

OpenStudy (anonymous):

if you are looking for the maclaurin series find the maclaurin series for \(\arctan(x)\) first then substitute

OpenStudy (anonymous):

The problem says to find the 9th derivative of \[f(x)=\tan^{-1}(\frac{x^{6}}{3}))\] so that at \[f^{9}(0)=?\]

OpenStudy (anonymous):

yikes really? you need an expression for the ninth derivative? the second derivative is hard enough

OpenStudy (anonymous):

I know, right? If I didn't know better, I'd swear he's singling me out. ;)

OpenStudy (anonymous):

I thought there might be a way to generalize the power series and get to it that way.

OpenStudy (anonymous):

i can't imagine how to do that sorry finding \(f^{9}(0)\) may not be that hard though

OpenStudy (anonymous):

Bummer. Yeah, I've seen typos in the homework before, but then again, many of the assigned problems have bordered on sadistic.

OpenStudy (anonymous):

i am fairly sure it is zero

OpenStudy (anonymous):

Yeah, I was, too, but Webwork didn't think so.

OpenStudy (anonymous):

but the maclaurin series will have only terms with powers of 6

OpenStudy (anonymous):

but for sure \(f^{9}(0)=0\)

OpenStudy (anonymous):

Actually, I get \[(\frac{x^3}{6})-[\frac{(\frac{x^3}{6})^{3}}{3}]-...\]

OpenStudy (anonymous):

Make that "+" on the end.

OpenStudy (thomas5267):

By the godlike Mathematica, it is \(f^9(0)=560\). Not sure how to get that though.

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

crap i had it backwards! it is \[\frac{x^3}{6}\] not \[\frac{x^6}{3}\] in which case there will be a degree 9 term

OpenStudy (anonymous):

Thomas5267, I've seen that somewhere else, as well, but not with any steps...though. Satellite73, I feel your pain. I do that all the time. :)

OpenStudy (anonymous):

So should I be taking the derivative of the second term?

OpenStudy (anonymous):

i get the idea i think

OpenStudy (anonymous):

find the maclaurin series for \[\arctan(\frac{x^3}{6})\] out to three terms the third term will have \(x^9\)

OpenStudy (anonymous):

the coefficient will be \[\frac{f^{9}(0)}{9!}\]

OpenStudy (anonymous):

Ah, I think that's it! Working through it now. *Brutal*.

OpenStudy (anonymous):

not that bad actually

OpenStudy (thomas5267):

I think this has something to do with this: \[ \sum_{n=0}^\infty x^n=\frac{1}{1-x} \]

OpenStudy (anonymous):

cheat the maclaurin series for arctangent is well known use it

OpenStudy (anonymous):

you can derive it using what @thomas5267 said but it is kind of annoying note that the derivative is \[\frac{1}{1+x^2}\] take \[\sum_{n=0}^\infty x^n=\frac{1}{1-x}\] replace \(x\) by \((-x^2)\) write the series integrate term by term too much work, look it up

OpenStudy (thomas5267):

Aren't we differentiating?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you find the maclaurin series for \[\frac{1}{1+x^2}\] then integrate term by term

OpenStudy (anonymous):

MacLaurin for arctan is where I got the x^9 in the first place. Arctan = x-(x^3/3)+(x^5/5)...

OpenStudy (anonymous):

in any case it is \[x-\frac{x^3}{3}+\frac{x^5}{5}-\frac{x^7}{7}+...\]

OpenStudy (anonymous):

replace \(x\) by \(\frac{x^3}{6}\)

OpenStudy (anonymous):

you get \(x^9\) in the second term

OpenStudy (anonymous):

if my arithmetic is correct you get \[\frac{x^9}{648}\] which means \[\frac{f^{9}(0)}{9!}=\frac{1}{648}\]

OpenStudy (thomas5267):

Weird. Mathematica returned -560.

OpenStudy (anonymous):

oooh an check it out!! \[\frac{9!}{648}=560\] just as mathematica said!!

OpenStudy (anonymous):

right i forgot the minus sign

OpenStudy (thomas5267):

Good job!

OpenStudy (anonymous):

that mathematica is very smart!!

OpenStudy (anonymous):

thnx trial an error strikes again

OpenStudy (thomas5267):

It is a software for mathematics, created by those guys behind Wolfram Alpha.

OpenStudy (anonymous):

yeah i used to use maple, now since i don't do any math i just use wolfram alpha on line

OpenStudy (anonymous):

Wow. Tough problem. So I want to be sure I understand here. You got the 648 from 3*216. The 9! is from it being the ninth term in the series, though the other powers don't appear. Is that right?

OpenStudy (anonymous):

I was so close, but I don't think I was raising 6^3.

OpenStudy (anonymous):

Thanks, both of you for your help!

OpenStudy (thomas5267):

You differentiate the second term of the series 9 times so you get 9!. \[ \frac{d^9}{dx^9}x^9=9! \] Then the other terms with powers of x vanish because the question is asking the value of \(f^9\) at \(x=0\).